40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10630     Entry time: Wed Oct 22 02:35:45 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Efforts at hopping PRMI to REFL165 

[EricQ, Jenne]

The first half of our evening was spent working on CARM and DARM in PRFPMI, and then we moved on to the PRMI part.

I moved the DARM ALSdiff -> TransDiff transition to be after the CARM ALScomm -> SqrtInvTrans transition in the carm_cm_up script.  After I did that, I succeeded every time (at least  10?  We did it many times) to get both CARM and DARM off of the ALS signals. 

We tried for a little while looking at transitioning to REFL11 normalized by the sum of the transmissions, but we kept losing lock.  We also several times lost lock at arm powers of a few, when we thought we weren't touching the IFO for any transitions.  Looking at the lockloss time series did not show any obvious oscillations in any of the _IN1 or _OUT channels for the length degrees of freedom, so we don't know why we lost lock, but it doesn't seem to be loop oscillations caused by changing optical gain.  Also, one time, I tried engaging Rana's "Lead 350" filter in FM7 of the CARM filter bank when we were on sqrtInvTrans for CARM, and the arm powers were around a few, but that caused the transmission signals to start to oscillate, and after one or two seconds we lost lock.  We haven't tried the phase lead filter again, nor have we tried the Boost2 that is in FM8. 

We increased the REFL11 analog gain from 0dB to 12dB, and then reset the dark offsets, but still weren't able to move CARM to normalized REFL11. Also, I changed the POP22 demod phase from 159 degrees to 139 degrees. This seems to be where the signal is maximized in the I-phase, while the arms are held off resonance, and also partway up the resonance peak. 

We then decided that we should go back to the PRMI situation before trying to reduce the CARM offset further.  We can robustly and quickly lock the PRMI on REFL33 while the arms are held off resonance with ALS.  So, we have been trying to acquire on  REFL33 I&Q, and then look at switching to REFL 165 I&Q.  It seems pretty easy to get PRCL over to REFL165 I (while leaving MICH on REFL33 I).  For REFL33, both matrix elements are +1.  For PRCL on REFL165, the matrix element is -0.08.  We have not successfully gotten MICH over to REFL 165 ever this evening. 

We went back and set the REFL165 I&Q offsets so that the outputs after the demod phase were both fluctuating around 0.  I don't know if they were around +/-100 because our dark offsets were bad or what, but we thought this would help.  We were still able to get PRCL transitioned no problem, but even after remeasuring the MICH REFL33 vs. REFL165 relative gains, we still can't transition MICH.  It seems like it's failing when the REFL33Q matrix element finally gets zeroed out, so we're not really getting enough signal in REFL165Q, or something like that, and throughout the rest of the transition we were depending entirely on REFL33Q. 

So. Plan:

  • Get PRMI on REFL165 while arms are held off resonance. 
    • May require PRCL-MICH FF decoupling, by combining error signals?
    • May require looking back at simulations to see what we expect the relative gains and signs to be.
  • Look at CARM loop stability in simulation for REFLDC, REFL11, and normalized REFL11.  Is there a stable loop path from about 100pm down to 0pm on normalized REFL11?
ELOG V3.1.3-